i have 18 coins that add up to $ 3.30 and all i had was quarters and dimes how many quarters did i have and how many dimes
Answer by Edwin McCravy(20086)


  q +   d =  18
25q + 10d = 330

Divide the 2nd equation through by 5

 q +  d = 18
5q + 2d = 66

Solve the first equation for d

d = 18 - q

Substitute (18 - q) for d

5q + 2(18 - q) = 66

5q + 36 - 2q = 66

3q = 30

 q = 10

Substitute in d = 18 - q

d = 18 - q = 18 - 10 = 8

10 quarters and 8 dimes

Edwin